Gas systems project

Europe/Zurich
155/R-047 (CERN)

155/R-047

CERN

10
Show room on map
Beatrice Mandelli (CERN), Frederic Merlet (CERN)

GAS M&O meeting 11 March 2019

Present: Beatrice, Fred, Patrick, Andrea, Michal, Roberto, LPh, Kacper, Luca

 

ALICE #1726

  • Preparing pipes for the TPC clean room —> to do this week or max the next
  • MTR: test still on-going. 
  • HMPID: safety valve mixer changed
  • MTR back-up: it has to be Freon —> LPh to check
  • MTR back-up is done with pressure reducer+rotameter: to check pressure when it opens 

 

ATLAS #1728

  • Starting some mechanics for RPC distribution racks 
  • Flowcell for new RPC racks —> Kacper to give the flow necessary —> to decide if 3 or 5 mm
  • ID flushing maintenance in w12/13

 

CMS #1724

  • RPC and GEM distribution racks almost completed —> negative side to send by end of the week
  • RPC distribution flowcell: verification cable and resistance for the rack in the workshop: file excel in Gas_Group/DistributionRack prepared by Patrick
  • Dummy chamber: to add the new flowcell to each manifold —> to start with racks in the surface —> Andrea to coordinate with Patrick
  • No quick on the manifold —> leak test postponed when the quick will arrive (so it will be done in the cavern)
  • Electrical chassis will be added at the end of the installation in the cavern
  • GEM pre-distribution and pump module in place —> missing pipes and cabling
  • Dummy chamber —> Beatrice too check with RPC people
  • DT ELMB reading card problem: no reading of the flow since 22th Jan (email from Daniel). DT ELMS Status: red square due to PowerSupply new logic —> it will be changed with new software
  • Valve island/pneumatic pipes —> need to stop RPC system for 2 weeks —> to coordinate with some standard maintenance 

 

LHCb #1727

  • RICH1 buffer ready —> waiting to have access in the cavern (priority to dismount detectors)
  • RICH2 batteries test —> done at 200 bar (valide for 140 bar)
  • RICH2 old compressor: asked to disconnect water and then dismount it —> this week
  • RICH2 compressor: to decide the way to recuperate gas —> Beatrice/Patrick to discuss
  • GEM stopped
  • SciFi: trolley ready, cabling done, leak test done. Missing software —> Michal this week
  • MWPC rack 64 there is a problem in the distribution rack: since mid January there were 200 l/h less —> Patrick to investigate with LHCb people

 

PIQUET

  • Nothing to report

 

GENERAL

  • Kacper to present the list of Krohne for Exhaust to check what we really need to change
  • Back-up with TESCOM —> to check where we have them in the back-up systems
  • Red pneumatic valves: several types of leaks and not spare pieces available on the market —> move to Swagelok when we need to change
  • Pressure sensor on filter for pump —> Kacper/Andrea to see. Two options: replace with the new inox (large surface) or keep the glass one

 

PLC/Software

  • Elog for CMS RPC PLC #1712
  • Profibus card: delay of maximum 3 weeks. In the meanwhile we use some cards from Marc
  • MCH PLC replacement/software: done but some bugs to fix
  • Next systems to do: CPV and TOTEM + debugging
  • 301 and 303 (on the mixer) are not supported for new WAGO  and there is problem with new software —> need to change —> Patrick
  • Excel file about CMS RPC distribution pressure sensor done by Patrick:  Gas_Group/CapteurDePression/2019 —> Beatrice/Patrick to check
There are minutes attached to this event. Show them.
    • 09:00 09:10
      Summary on first PLC replacement and crash of CMS RPC PLC 10m
      Speaker: Michal Zbigniew Zimny (CERN)
    • 09:10 09:15
      Piquet interventions 5m

      http://eloggas.cern.ch/GAS_PIQUET/

      Speakers: Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Frederic Merlet (CERN), Louis-Philippe De Menezes (CERN), Patrick Carrie (CERN), Kacper Jerzy Kapusniak (CERN), Michal Zbigniew Zimny (CERN)
    • 09:15 09:20
      Activities on-going 5m
      Speakers: Beatrice Mandelli (CERN), Frederic Merlet (CERN)
    • 09:20 09:40
      Plan of LS2 activities 20m
      Speakers: Beatrice Mandelli (CERN), Frederic Merlet (CERN), Roberto Guida (CERN)
    • 09:40 11:05
      Gas systems for the LHC experiments
    • 11:05 11:15
      FSU & Jobs status 10m
      Speaker: Frederic Merlet (CERN)
    • 11:15 11:25
      AoB 10m
      Speakers: Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Frederic Merlet (CERN), Louis-Philippe De Menezes (CERN), Patrick Carrie (CERN), Roberto Guida (CERN)
Your browser is out of date!

If you are using Internet Explorer, please use Firefox, Chrome or Edge instead.

Otherwise, please update your browser to the latest version to use Indico without problems.

×